(10 marks) Suppose you have three production facilities that are polluting a river. Each emits 10 units of pollution. Their marginal cost functions for reducing emissions are, respectively, ??1 = $3, ??2 = $4, and ??3 = $5. (a) If the objective is to cut emissions in half (to 15) cost-effectively, how much reduction should be assigned to each firm?

(b) What would be the total variable cost of controlling these emissions?

(c) What would be the total variable cost that would result from forcing each facility to control one-half of its emissions? Is this different from the cost associated with the cost-effective allocation? Why or why not?
Answer #1

a) as firm 1 is lowest MC firm, so firm 1 should abate all 10 units, & remaining 5 units are abated by second firm, as it's MAC is second lowest

firm 1 should cut emissions by 10

Firm 2 should cut emissions by 5

b) TVC = 10*3 + 5*4

= 30+20

= $ 50

c) if each abates by 5

Then TAC = 5*3 + 5*4 + 5*5

= 15+20+25

= $ 60

so yes, this is different & higher than in part a) bcoz it does involve the cost minimization

the firm with lower marginal abatement cost MAC should Abate more pollution units

​​​​​